3. Using the Harris-Benedict equation, calculate the caloric requirement for a 45 year old women, who is confined to a bed and just came out of surgery. (weight: 121 lb & height: 5 ft 3 in.)
Answer #1

Women: BMR = 655 + ( 4.35 x weight in pounds ) + ( 4.7 x height in inches ) - ( 4.7 x age in years )

4.35 x weight in pounds

= 4.35 x 121 lb

= 526.35 pounds

( 4.7 x height in inches )

4.7 x 63 inches

= 296.1

( 4.7 x age in years)

4.7 x 45

= 211.5

= 655 + 526.35 pounds + 296.1 inches - 211.5 yrs

= 1477.45- 211.5

= 1265.95

Calories needed each day Sedentary (little or no exercise): BMR x 1.2

= 1265.95 x 1.2

= 1519.14